# No-Rank Tensor Decomposition Using Metric Learning

### v1

@@ 5.2. Metric Learning Performance on Brain Connectivity Data
To evaluate metric learning framework on a complex, high-dimensional biomedical problem, we applied it to the Autism Brain Imaging Data Exchange (ABIDE) [di2017enhancing] dataset. This dataset contains resting-state functional MRI (rs-fMRI) data from 871 subjects (403 with Autism Spectrum Disorder, ASD, and 468 typically developing controls). Each subject is represented by a $111\times 111$ functional connectivity matrix, which are flattened and normalized to serve as input. The clinical labels (ASD vs. control) provide the semantic similarity relationships for guiding the metric learning process.
@@
The clustering results, presented in Table 3, demonstrate that metric learning framework significantly outperforms all baseline methods. It achieves a near-perfect Silhouette Score of 0.9880, indicating exceptional separation between the ASD and control groups. This represents a substantial improvement over the best unsupervised method, UMAP (0.6194). The superior cluster compactness and separation are further confirmed by the lowest Davies-Bouldin Index (0.0186) and an exceptionally high Calinski-Harabasz Index (31912.94), which is two orders of magnitude greater than UMAP’s score (311.35). Crucially, the proposed method is the only one to achieve meaningful alignment with clinical labels, as evidenced by its strong external validation metrics (ARI: 0.3002, NMI: 0.2372), while all other methods yield scores near zero, failing to recover the underlying diagnostic structure. This performance highlights a key limitation of fixed-rank tensor decompositions: both CP and Tucker variants show severe sensitivity to rank selection and achieve negligible clustering quality and external validation scores across all tested ranks (CP: 0.0022-0.0070, Tucker: 0.0046-0.0061 Silhouette; ARI/NMI ¡ 0.02). In contrast, the no-rank approach avoids this hyperparameter sensitivity altogether while discovering semantically meaningful structure.
@@ 5.3. Metric Learning Performance on Simulated Datasets
Figure 4. Example images from simulated datasets. Top: Galaxy morphology classes (Elliptical, Spiral, Lenticular, Irregular). Bottom: Crystal structure classes (Cubic, Hexagonal, Tetragonal, Orthorhombic).
@@
Metric learning achieves near-perfect Silhouette scores (0.9999) and optimal Davies-Bouldin indices (0.0001) on both datasets, indicating ideal cluster separation and compactness. The dramatic improvements in Calinski-Harabasz scores—over seven orders of magnitude higher than PCA—and the significantly enhanced Separation Ratios demonstrate the method’s exceptional ability to create a discriminative embedding space.
This performance comes with the expected trade-off: metric learning exhibits moderately lower Continuity and Trustworthiness scores compared to some unsupervised methods. This indicates that while global cluster structure is optimized for class separation, some local neighborhood relationships from the original pixel space are distorted—a deliberate consequence of reorganizing the space around semantic class identity rather than pixel-level similarity.
@@
The quantitative superiority is visually confirmed in Figure 5. The metric-learned embeddings form tight, well-separated clusters with minimal overlap, contrasting with the more entangled structures from unsupervised techniques. This clear visual separation underscores the advantage of learning a task-specific distance metric that directly optimizes for class discrimination.
Figure 5. Embedding visualizations. Two Top Rows Galaxy morphology dataset. Two Bottom Rows Crystal structure dataset. Metric learning produces the most distinct and compact clusters, demonstrating its effectiveness in learning semantically meaningful representations.
@@ 5.4. Comparison with Tensor Decomposition Methods in Terms of Reconstruction
In addition to the metric learning approach, we compare against two fundamental tensor decomposition methods: CP and Tucker decomposition. For a third-order tensor $\mathcal{X}\in\mathbb{R}^{I\times J\times K}$ representing our scientific dataset (samples $\times$ height $\times$ width), these decompositions are defined as follows:
@@ 5.4.1. Comparative Analysis Metrics
The fundamental difference lies in how each method handles dimensionality. While tensor decomposition methods require explicit rank specification RR—typically fixed unless an adaptive approach is implemented [DektorRodgersVenturi2021, sedighin2021adaptive]—they also necessitate rank optimization and multiple rank trials, often guided by theoretical rank estimation. In contrast, the metric learning approach operates with an implicit latent dimension $d$, features an adaptive hierarchical representation, and requires only a single training procedure. This leads to a fully data-driven representation learning process.
A comparison between the metric learning approach and baseline tensor decomposition methods in terms of reconstruction error and explained variance over the same datasets discussed above is provided in Table 5. The results for CP and Tucker are reported for best-rank approximation for $R=2,3,5,10,15,20$.
@@
Dataset Method Reconstruction Error Explained Variance LFW CP Decomposition 0.5300 ($R=20$) 0.7191 Tucker Decomposition 0.4908 ($R=(20,20,20)$) 0.7591 Metric Learning 0.0991 0.9901 Olivetti CP Decomposition 0.6441 ($R=5$) 0.5852 Tucker Decomposition 0.4326 ($R=(20,20,20)$) 0.8129 Metric Learning 0.1001 0.9899 ABIDE CP Decomposition 0.6070 ($R=20$) 0.6315 Tucker Decomposition 0.4577 ($R=(20,20,20)$) 0.7905 Metric Learning 0.0139 0.9998 Galaxy Morph. CP Decomposition 0.5049 ($R=10$) 0.7451 Tucker Decomposition 0.4855 ($R=(20,20,20)$) 0.7643 Metric Learning 0.0685 0.9953 Crystal Struc. CP Decomposition 0.4088 ($R=5$) 0.8329 Tucker Decomposition 0.3819 ($R=(5,5,5)$) 0.8542 Metric Learning 0.0782 0.9938
The reconstruction visualization demonstrates the process of approximating original tensors from their compressed representations obtained through various decomposition methods. For a third-order tensor $\mathcal{X}\in\mathbb{R}^{I\times J\times K}$, each method employs distinct reconstruction mechanisms. In CP decomposition, the original tensor is reconstructed from rank-1 components through the summation $\mathcal{\hat{X}}=\sum_{r=1}^{R}\lambda_{r}\cdot\mathbf{a}_{r}\circ\mathbf{b}_{r}\circ\mathbf{c}_{r}$, where $\lambda_{r}$ represents scaling weights and $\mathbf{a}_{r},\mathbf{b}_{r},\mathbf{c}_{r}$ are factor matrices capturing different modes of variation. The Tucker decomposition utilizes a more flexible reconstruction via $\mathcal{\hat{X}}=\mathcal{G}\times_{1}\mathbf{A}\times_{2}\mathbf{B}\times_{3}\mathbf{C}$, where $\mathcal{G}$ is the core tensor encoding interactions between factors and $\times_{n}$ denotes the n-mode product. For metric learning, reconstruction is achieved through linear regression $\mathbf{\hat{X}}_{\text{flat}}=\mathbf{W}\mathbf{Z}+\mathbf{b}$, where $\mathbf{Z}$ represents the learned embeddings and $\mathbf{W}$ maps these back to the original space. The visual comparison reveals that tensor decompositions excel at structural preservation due to their explicit reconstruction formulas, while metric learning prioritizes discriminative feature retention over perfect reconstruction fidelity, reflecting their different optimization objectives in capturing brain connectivity patterns. The constructed faces for Olivetti dataset and cluster visualizations for LFW datasets are shown in Figure 6 and 7, respectively.
@@
Figure 6. Visual comparison of original tensor from Olivetti face dataset and their reconstructions using different decomposition methods. Here, the results for metric learnign are derived with only 50 epochs.
@@
Figure 7. Two-dimensional manifold visualizations of face embeddings from the LFW dataset generated by different dimensionality reduction and tensor decomposition methods. Each subplot shows the embedding space where points represent individual face images colored by identity, demonstrating the clustering performance and separability achieved by each method.
@@
Figure 8 demonstrates the two-dimensional manifold visualizations of brain embeddings, comparing original data with CP decomposition, Tucker decomposition, and metric learning approaches. All methods successfully preserve the symmetric structure and hub connectivity patterns essential for neurological analysis.
@@
Figure 8. Two-dimensional manifold visualizations of brain embeddings from the ABIDE dataset generated by different dimensionality reduction and tensor decomposition methods. Each subplot shows the embedding space where points represent individual face images colored by identity, demonstrating the clustering performance and separability achieved by each method. The best reported model accuracy for CP with $R=10$ is 0.6000, for Tucker decompostion with $R=(20,20,20)$ is 0.7000 and for Metric learning is 0.9900.
@@ 5.4.2. Empirical Rank Robustness
Our experiments demonstrate that while tensor decomposition performance varies significantly with rank choice, the metric learning approach maintains consistent performance across different dataset complexities. This robustness stems from non-linear transformations that can capture complex interactions without explicit high-rank decomposition, hierarchical feature learning through convolutional layers that automatically organize features by complexity and task-oriented optimization where the representation is learned specifically for the scientific classification task.
@@
This section presents a comprehensive comparative analysis between the proposed Metric Learning method and Transformer-based models across five diverse datasets spanning brain connectivity, face recognition, and scientific image classification. The objective is not to claim universal superiority but to delineate the specific scenarios, particularly data-constrained environments, where our method offers distinct advantages over Transformer architectures [vaswani2017attention]. Experiments were conducted on the same datasets as above, namely ABIDE ($111\times 111$ connectivity matrices), LFW faces ($50\times 37$), Olivetti faces ($64\times 64$), Galaxy morphology ($64\times 64$), and Crystal structure ($64\times 64$), demonstrating the generalizability of our findings across multiple domains.
@@ 5.5.2. Advantages of the Proposed Metric Learning Framework
The proposed Metric Learning method demonstrated consistent operational success and competitive performance across all datasets, through geometric learning principles. The proposde method executed successfully on all datasets across all experimental conditions. The most significant advantage emerged in small-data regimes where Transformers remained inapplicable.
@@ 6. Conclusion
6.1. Advantages of the Metric Learning Framework
The metric learning framework demonstrates consistent superiority over traditional dimensionality reduction techniques, including tensor decomposition methods like CP and Tucker. The performance advantages stem from fundamental differences in optimization objectives, architectural flexibility, and learning capabilities.
Discriminative vs. Reconstructive Optimization: Traditional methods like PCA, CP, and Tucker decompositions focus on reconstructive objectives—preserving maximum information for data reconstruction. In contrast, the proposed framework employs discriminative optimization through triplet losses that explicitly maximize inter-class distances while minimizing intra-class variations. This direct alignment with clustering objectives yields embedding spaces where class separation is the primary goal, evidenced by superior Adjusted Rand Index (ARI) scores across all evaluated datasets.
Adaptive Architecture and Rank Flexibility: Tensor decomposition methods are constrained by pre-defined rank parameters that limit their adaptability to complex data structures. This framework operates without such constraints, with the embedding dimension serving as a flexible hyperparameter rather than a structural limitation. This adaptability is demonstrated in rank sensitivity analysis, where metric learning maintained consistent high performance while CP and Tucker methods showed significant variation across different ranks.
Integrated Enhancement Strategies: The framework’s modular architecture supports multiple performance-enhancing strategies unavailable to fixed decomposition algorithms. Data augmentation improves robustness to biological variability, autoencoder pre-training accelerates convergence, ensemble methods enhance stability, and optional UMAP post-processing refines visualizations. This integrative capability allows the framework to adapt to diverse data characteristics and computational requirements.
Structural Preservation and Efficiency: Beyond superior clustering performance, proposed approach demonstrates better structural preservation through higher trustworthiness metrics and separation ratios. The end-to-end deep architecture enables hierarchical feature learning, capturing multi-scale patterns essential for complex data. Additionally, the trained model offers efficient inference through simple forward passes, making it suitable for real-time applications where repeated tensor decompositions would be computationally prohibitive.
These advantages—discriminative optimization, architectural flexibility, integrative capabilities, and computational efficiency—collectively position metric learning framework as a more powerful paradigm for applications where semantic similarity and clinical discriminability are paramount concerns.
6.2. Limitations and Future Work
@@
7. Data Statement
@@
    Publicly available Labeled Faces in the Wild (LFW) dataset [huang2008labeled].
@@
    Publicly available Olivetti Faces dataset [pedregosa2011scikit].
@@
    Publicly available Autism Brain Imaging Data Exchange (ABIDE) dataset [di2017enhancing].

### v2

@@ 5.2. Metric Learning Performance on Brain Connectivity Data
Unlike unsupervised baselines, the metric learning model leverages diagnostic labels during training to shape the embedding geometry. To evaluate metric learning framework on a complex, high-dimensional biomedical problem, we applied it to the Autism Brain Imaging Data Exchange (ABIDE) [12] dataset. This dataset contains resting-state functional MRI (rs-fMRI) data from 871 subjects (403 with Autism Spectrum Disorder, ASD, and 468 typically developing controls). Each subject is represented by a $111\times 111$ functional connectivity matrix, which are flattened and normalized to serve as input. The clinical labels (ASD vs. control) provide the semantic similarity relationships for guiding the metric learning process.
@@
The clustering results in Table 3 demonstrate that the proposed metric learning framework yields substantially more discriminative embeddings than all baseline methods. Owing to its supervised objective, the learned embedding space exhibits extremely tight intra-class compactness and strong inter-class separation, reflected by a very high Silhouette score (0.9932) and a low Davies–Bouldin index (0.0186). The large Calinski–Harabasz value further indicates strong between-group separation relative to within-group dispersion, which is expected when semantic labels explicitly guide the geometry of the embedding space.
In contrast, unsupervised baselines—including UMAP, t-SNE, and all fixed-rank tensor decompositions—optimize geometric or reconstruction-based objectives without access to diagnostic labels. As a result, while some methods (e.g., UMAP) yield visually compact clusters, they fail to align with clinical structure, as evidenced by near-zero ARI and NMI scores. This discrepancy highlights a fundamental limitation of reconstruction-driven and rank-constrained tensor methods in clinical settings, where discriminative structure does not necessarily correspond to low-rank variance.
Notably, CP, Tucker, and t-SVD decompositions exhibit strong sensitivity to rank selection and consistently produce negligible external validation scores across all tested ranks (Silhouette ¡ 0.01, ARI/NMI ¡ 0.02). By contrast, the proposed no-rank metric learning approach avoids explicit rank tuning and directly optimizes clinically meaningful similarity, resulting in substantially improved alignment with diagnostic labels (ARI = 0.3002, NMI = 0.2372). These results suggest that discriminative metric learning provides a more appropriate inductive bias for brain connectivity analysis than reconstruction-based tensor factorization.
@@ 5.3. Metric Learning Performance on Simulated Datasets
Figure 5. Galaxy morphology & Crystal structure:Example images from simulated datasets. Top: Galaxy morphology classes (Elliptical, Spiral, Lenticular, Irregular). Bottom: Crystal structure classes (Cubic, Hexagonal, Tetragonal, Orthorhombic).
@@
Metric learning achieves near-perfect Silhouette scores ($\approx$1.0) and near-zero Davies–Bouldin indices on both simulated datasets, reflecting almost complete separation between classes. This behavior is expected in a supervised metric learning setting when class boundaries are well-defined by construction. The extremely large Calinski–Harabasz values and Separation Ratios arise from the explicit maximization of inter-class margins and contraction of intra-class distances, rather than from variance-based or reconstruction-driven objectives optimized by unsupervised baselines.
This performance comes with the expected trade-off: metric learning exhibits moderately lower Continuity and Trustworthiness scores compared to some unsupervised methods. This indicates that while global cluster structure is optimized for class separation, some local neighborhood relationships from the original pixel space are intentionally distorted—a known and deliberate consequence of prioritizing semantic class identity over raw pixel similarity.
@@
The quantitative superiority is visually confirmed in Figures 16 and 6 (see Appendix 7). The metric-learned embeddings form tight, well-separated clusters with minimal overlap, contrasting with the more entangled structures from unsupervised techniques. This clear visual separation underscores the advantage of learning a task-specific distance metric that directly optimizes for class discrimination. Visual comparison of original tensors from Galaxy morphology and Crysyal Structure datasets and their reconstructions using deep laerning models, VAE and DEC are shown in Figure 17 and 7 (see Appendix 7), respectively.
Figure 6. Crystal structure: Embedding visualizations for Crystal structure dataset: Metric learning produces the most distinct and compact clusters, demonstrating its effectiveness in learning semantically meaningful representations (see also Figure 16).
Figure 7. Galaxy morphology : Visual comparison of original tensor from Galaxy morphology dataset and their reconstructions using deep laerning models, VAE and DEC. All models recived similar pararmerts, number of epochs, normalization and are evaluated on the same metrics (see also Figure 17).
@@ 5.4. Comparison with Tensor Decomposition Methods in Terms of Reconstruction
In addition to the metric learning approach, we compare against two fundamental tensor decomposition methods: CP, Tucker and t-SVD decompositions, as well deep learning models VAE and DEC. For a third-order tensor $\mathcal{X}\in\mathbb{R}^{I\times J\times K}$ representing our scientific dataset (samples $\times$ height $\times$ width), these decompositions are defined as follows:
@@ 5.4.1. Comparative Analysis Metrics
The fundamental difference lies in how each method handles dimensionality. While tensor decomposition methods require explicit rank specification RR—typically fixed unless an adaptive approach is implemented [11, 41]—they also necessitate rank optimization and multiple rank trials, often guided by theoretical rank estimation. In contrast, the metric learning approach operates with an implicit latent dimension $d$, features an adaptive hierarchical representation, and requires only a single training procedure. This leads to a fully data-driven representation learning process.
A comparison between the metric learning approach and baseline tensor decomposition methods in terms of reconstruction error and explained variance over the same datasets discussed above is provided in Table 5. The results for CP, Tucker and t-SVD are reported for best-rank approximation for $R=2,3,5,10,15,20$.
@@
Dataset Method Reconstruction Error Explained Variance LFW CP Decomposition 0.5300 ($R=20$) 0.7191 Tucker Decomposition 0.4908 ($R=(20,20,20)$) 0.7591 t-SVD 0.4908 ($R=20$) 0.7590 VAE 0.6804 0.4153 DEC 0.4835 0.5165 Metric Learning 0.0991 0.9901 Olivetti CP Decomposition 0.6441 ($R=5$) 0.5852 Tucker Decomposition 0.4326 ($R=(20,20,20)$) 0.8129 t-SVD 0.4326 ($R=20$) 0.8129 VAE 0.6753 0.4190 DEC 0.4087 0.5913 Metric Learning 0.1001 0.9899 ABIDE CP Decomposition 0.6070 ($R=20$) 0.6315 Tucker Decomposition 0.4577 ($R=(20,20,20)$) 0.7905 t-SVD 0.4577 ($R=20$) 0.7905 VAE 0.8918 0.1122 DEC 0.5709 0.4291 Metric Learning 0.0139 0.9998 Galaxy Morph. CP Decomposition 0.5049 ($R=10$) 0.7451 Tucker Decomposition 0.4855 ($R=(20,20,20)$) 0.7643 t-SVD 0.7190($R=20$) 0.4832 VAE 0.5572 0.4895 DEC 0.2886 0.7114 Metric Learning 0.0685 0.9953 Crystal Struc. CP Decomposition 0.4088 ($R=5$) 0.8329 Tucker Decomposition 0.3819 ($R=(5,5,5)$) 0.8542 t-SVD 0.4440 ($R=5$) 0.8026 VAE 0.8705 0.1532 DEC 0.1851 0.8149 Metric Learning 0.0782 0.9938
The reconstruction visualization demonstrates the process of approximating original tensors from their compressed representations obtained through various decomposition methods. For a third-order tensor $\mathcal{X}\in\mathbb{R}^{I\times J\times K}$, each method employs distinct reconstruction mechanisms.
In CP decomposition, the original tensor is reconstructed from rank-1 components through the summation $\mathcal{\hat{X}}=\sum_{r=1}^{R}\lambda_{r}\cdot\mathbf{a}_{r}\circ\mathbf{b}_{r}\circ\mathbf{c}_{r},$ where $\lambda_{r}$ represents scaling weights and $\mathbf{a}_{r},\mathbf{b}_{r},\mathbf{c}_{r}$ are factor matrices capturing different modes of variation.
The Tucker decomposition utilizes a more flexible reconstruction via $\mathcal{\hat{X}}=\mathcal{G}\times_{1}\mathbf{A}\times_{2}\mathbf{B}\times_{3}\mathbf{C},$ where $\mathcal{G}$ is the core tensor encoding interactions between factors and $\times_{n}$ denotes the $n$-mode product.
The t-SVD framework reconstructs the tensor through the tensor product $\mathcal{\hat{X}}=\mathcal{U}*\mathcal{S}*\mathcal{V}^{\top},$ where * denotes the t-product defined in the Fourier domain, and $\mathcal{S}$ is a f-diagonal tensor containing singular values across frontal slices. This formulation preserves multi-way correlations along the third mode and yields an optimal low-rank approximation under the tensor nuclear norm. For metric learning, reconstruction is achieved through linear regression $\mathbf{\hat{X}}_{\text{flat}}=\mathbf{W}\mathbf{Z}+\mathbf{b},$ where $\mathbf{Z}$ represents the learned embeddings and $\mathbf{W}$ maps these back to the original space.
For the deep learning models, reconstruction is achieved via their respective decoders. For VAE the encoder maps the input tensor into a latent Gaussian distribution, and the decoder reconstructs $\hat{\mathcal{X}}$ by sampling from this distribution and applying non-linear transformations through the network layers. For DEC the encoder produces a latent embedding which is optimized to match a target distribution for clustering, and the reconstruction $\hat{\mathcal{X}}$ is obtained by passing these embeddings through a decoder network trained to minimize reconstruction loss (e.g., mean squared error) while aligning cluster assignments. These decoders allow VAE and DEC to approximate the original tensor from learned latent representations, highlighting the difference between reconstruction-focused methods and our metric learning approach, which prioritizes discriminative embeddings over exact reconstruction.
The visual comparison reveals that tensor decomposition methods (CP, Tucker, and t-SVD) emphasize structural preservation through explicit algebraic reconstruction, whereas metric learning prioritizes discriminative feature retention over exact reconstruction fidelity. This distinction reflects their different optimization objectives in capturing brain connectivity patterns. The constructed faces for the Olivetti dataset and cluster visualizations for the LFW dataset are shown in Figure 8 and Figure 10, respectively. Figure 9 shows the functional matrix reconstruction for ABIDE dataset, compared with VAE, and DEC methods. For other compariosn see also Figures 1417 in Appendix 7).
@@
Figure 8. Olivetti Faces : Visual comparison of original tensor from Olivetti face dataset and their reconstructions using different decomposition methods. Here, the results for metric learnign are derived with 50 epochs (see also Figure 14 in Appendix 7).
Figure 9. ABIDE dataset : Visual comparison of original tensor from functional connectivity matrices of ABIDE dataset and their reconstructions using deep laerning models, VAE and DEC. All models recived similar pararmerts, number of epochs, normalization and are evaluated on the same metrics.
@@
Figure 10. LFW Dataset : Two-dimensional manifold visualizations of face embeddings from the LFW dataset generated by different dimensionality reduction and tensor decomposition methods. Each subplot shows the embedding space where points represent individual face images colored by identity, demonstrating the clustering performance and separability achieved by each method (similar figure for Olivetti Faces dataset is added to Appendix 7, Figure 13).
@@
Figure 11 demonstrates the two-dimensional manifold visualizations of brain embeddings, comparing original data with CP decomposition, Tucker and t-SVD decompositions, and metric learning approaches. All methods successfully preserve the symmetric structure and hub connectivity patterns essential for neurological analysis.
@@
Figure 11. ABIDE dataset : Two-dimensional manifold visualizations of brain embeddings from the ABIDE dataset generated by different dimensionality reduction and tensor decomposition methods. Each subplot shows the embedding space where points represent individual face images colored by identity, demonstrating the clustering performance and separability achieved by each method. The best reported model accuracy for CP with $R=10$ is 0.6000, for Tucker and t-SVD decompostion with $R=(20,20,20)$ and $R=20$ respectively, is around 0.7000, and for Metric learning is 0.9900.
@@ 5.4.2. Empirical Rank Robustness
Our experiments demonstrate that while tensor decomposition performance varies with rank choice, the metric learning approach maintains consistent performance across different dataset complexities. This robustness stems from non-linear transformations that can capture complex interactions without explicit high-rank decomposition, hierarchical feature learning through convolutional layers that automatically organize features by complexity and task-oriented optimization where the representation is learned specifically for the scientific classification task.
@@
This section presents a comprehensive comparative analysis between the proposed Metric Learning method and Transformer-based models across five diverse datasets spanning brain connectivity, face recognition, and scientific image classification. The objective is not to claim universal superiority but to delineate the specific scenarios, particularly data-constrained environments, where our method offers distinct advantages over Transformer architectures [47]. Experiments were conducted on the same datasets as above, namely ABIDE ($111\times 111$ connectivity matrices), LFW faces ($50\times 37$), Olivetti faces ($64\times 64$), Galaxy morphology ($64\times 64$), and Crystal structure ($64\times 64$), demonstrating the generalizability of our findings across multiple domains.
@@ 5.5.2. Advantages of the Proposed Metric Learning Framework
The proposed Metric Learning method demonstrated consistent operational success and competitive performance across all datasets, through geometric learning principles. The proposde method executed successfully on all datasets across all experimental conditions. The most advantage emerged in small-data regimes where Transformers remained inapplicable.
@@ 6. Conclusion
Advantages of the Metric Learning Framework
The proposed metric learning framework demonstrates strong performance for tasks that require semantically meaningful embeddings, such as clustering and representation analysis. In contrast to classical dimensionality reduction techniques—including CP, Tucker, and t-SVD—which primarily optimize reconstruction error under fixed-rank constraints, our approach explicitly optimizes similarity structure through task-driven loss functions. This design yields representations that are particularly well-suited for downstream tasks involving discrimination, retrieval, and semantic organization of data.
The flexibility of the framework avoids explicit rank selection, allowing the embedding dimension to adapt to data complexity, while modular strategies such as data augmentation, pre-training, and ensemble modeling further improve robustness and generalization. Although the deep network provides substantial modeling capacity, the observed performance improvements are largely attributable to the carefully designed metric learning objectives that directly encode semantic structure. Overall, these properties make the proposed framework a practical and adaptable tool for applications in which meaningful representation learning is prioritized over exact signal reconstruction.
Future Work
@@
Code Availability & Data Statement
An implementation of the proposed method is available at https://github.com/mbagherian/No-Rank-Tensor-Decomposition.
@@
    Publicly available Labeled Faces in the Wild (LFW) dataset https://www.kaggle.com/datasets/jessicali9530/lfw-dataset [22].
@@
    Publicly available Olivetti Faces dataset (credit to AT&T Laboratories Cambridge http://www.cl.cam.ac.uk/research/dtg/attarchive/facedatabase.html and [37])
@@
    Publicly available Autism Brain Imaging Data Exchange (ABIDE) dataset https://fcon_1000.projects.nitrc.org/indi/abide/ [12].
@@
Author Contribution Declarations
M.B. developed the research idea, performed all analyses and experiments, and prepared the manuscript.
Funding
This work was supported by the startup fund (Grant number ASE016) provided by the College of Science and Engineering (CoSE) at Idaho State University.